OpenStudy (anonymous):

Find the x-intercepts of the parabola with vertex (-3,16) and y-intercept (0,-20). Write your answer in this form: (x1,y1),(x2,y2). If necessary, round to the nearest hundredth. Use the method of completing the square to write the equation of the given parabola in this form: (y – k) = a(x – h)2 where a ≠ 0, (h,k) is the vertex, and x = h is the axis of symmetry

OpenStudy (campbell_st):

substitute the vertex into the equation \[(y - 16) = a(x + 3)^2\] to find the value of a, substitute the y-intercept x = 0, y = -20 into the equation and solve for a. this will give the actual equation of the parabola... it can be re-written as \[y = a(x + 3)^2 + 16\] distribute and collect like terms... then you will have the parabola in the standard form. Use the general quadratic formula to find x intercepts